OpenStudy (anonymous):

simplify 4√ 6 / √ 30 by rationalizing the denominator?

OpenStudy (anonymous):

\[\frac{ 4\sqrt{6} }{ \sqrt{30}}\]

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):